HETAV PATEL

Grade 11,

the reaction A->B follows a first order kinetics. the time taken for 0.8 mole of A to produce 0.6 mole of B is 1 hour . what is the time taken for conversion of 0.9 mole of A to produce 0.675 mole of B? plzz tell how did u find the answer

the reaction A->B follows a first order kinetics. the time taken for 0.8 mole of A to produce 0.6 mole of B is 1 hour . what is the time taken for conversion of 0.9 mole of A to produce 0.675 mole of B? plzz tell how did u find the answer

Grade:12

3 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

The integrated rate law for a 1st order reaction is:

ln([A]0) - ln([A]) = kt

One interesting thing about first order reactions is that the time required for 1/2 of the beginning reagent to react is independant of the conentration of the reagent. This can easily be seen by combining the two logs in the above integrated rate expression

ln([A]0/[A]) = kt

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

vikas askiitian expert
13 years ago

for first order reaction ,

t = 2.303log(a/a-x) / k                    ....................1

k is rate constant ..

when .8 mol of A produces .6mol of B then .8-.6 = .2 mol of A is left ...

final moles A = a-x = .2

initial moles = a = .8

time taken = 1hr

by plugging these values in eq 1 , we get

k = 2.303log4 ...............2

 

in second case , .9mol of A produces .675mols of B so remaining moles

of A are .9-.675 = .225

final moles of A = a-x = .225

initial mole = .9

now againg using eq 1

t = 2.303log(.9/.225)/k

  =2.303log4/k

k = 2.303log4 , so

t = 1hr

 

therefore time in second part will be same 1 hr...

 

appr0ve if u like this ans

ankit singh
3 years ago
or first order reaction ,
t = 2.303log(a/a-x) / k ....................1
k is rate constant ..
when .8 mol of A produces .6mol of B then .8-.6 = .2 mol of A is left ...
final moles A = a-x = .2
initial moles = a = .8
time taken = 1hr
by plugging these values in eq 1 , we get
k = 2.303log4 ...............2
in second case , .9mol of A produces .675mols of B so remaining moles
of A are .9-.675 = .225
final moles of A = a-x = .225
initial mole = .9
now againg using eq 1
t = 2.303log(.9/.225)/k
=2.303log4/k
k = 2.303log4 , so
t = 1hr
therefore time in second part will be same 1 hr..
Hope it will help you...
Pls mark me brainliest...